QUESTION#1

National dental clinic specializes in root canal operation without administration of pain-killing drugs. If the output (Q) is measured as number of root canals performed on daily basis, define the short run measures of costs: FC, VC,TC,MC,AFC,AVC,ATC Do the necessary calculations and fill the spaces in the table given below:(Copy & Complete the table below in your answer sheet only) [5]
Q FC VC TC MC AFC AVC ATC
1 $13 $38
2 $28
3 $70
4 $64
5 $22
6 $108
7 $133
8 $20













Question # 2 Answer the following questions. (Explain graphically and assume all necessary figures to support your answer where necessary) [05]
I. Explain graphically how indifference curve analysis can be used to derive a demand curve?

II. Explain the Law of diminishing return and why is it applicable especially in agriculture sector?
1
Expert's answer
2021-08-16T17:47:19-0400

QUESTION 1


FC is that part of TC which is indepenent of Q.

VC is that part of TC which is depenent on Q.

TC is the sum of VC and FC.

MC is the addition to TC due to production of an additional unit.

AFC is the per unit fixed cost.

AVC is the per unit variable cost.

ATC is the per unit total cost.


The complete table




"FC = TC - VC = 38 - 13 = 25." It is same for all levels of Q.

"VC = AVC\times Q" or "VC = TC - FC"

"TC = ATC\times Q" or "TC = VC + FC"

MC for Ist unit = TC - FC; For other units = Change in TC

"AFC = \frac{FC}{Q}"

"AVC = \frac{VC}{Q}"

"ATC = \frac{TC}{Q}"



QUESTION 2


I

the demand curve slope downward to the right, because as the price of a good falls both the substitution effect and income effect pull together in increasing the quantity demanded of the good.




From the above diagram, the demand curve DD is derived from the indifference curve in the top panel whereby, the X-axis shows the quantity demanded as in indifference curve and Y-axis is the price per unit of good X. In order to obtain the demand curve, various points K, L, S and T representing the demand schedule are plotted. By joining the points K, L, Sand T we get the required demand curve DD.


II

The law of diminishing returns refers to the extra returns obtained by employing one more unit of variable input which decreases as a firm keeps on increasing the variable input without increasing the fixed input.The law is applied in agriculture because land is a fixed input in the short run.